# Accuracy and Political Bias of News Source Credibility Ratings by Large Language Models

### v2

@@ Results
Figure 3: Heatmap of source credibility rating correlation (Spearman) among different LLMs and human experts. Results in the upper right triangle of the heatmap are based on 3,606 (48.0%) sources rated by all LLMs. Results in the lower left triangle are based on the sources rated by both raters in comparison.
@@
Let us measure how well the LLM ratings correlate with each other and align with those from human experts. For each pair of raters (LLMs or human experts), we calculate the Spearman rank correlation coefficient $\rho$ between their ratings. Given that some LLMs only provide ratings for a subset of the sources, we use two different approaches to measure the correlations. First, we include only the 3,606 (48.0%) sources that are rated by all LLMs. Second, for each pair of raters, we focus on their intersection. The results from both approaches are shown in Figure 3 and demonstrate similar patterns. Therefore, our discussion below will focus on the results from the first approach.
@@
All correlation coefficients in Figure 3 are positive and statistically significant with $p<0.001$. We find a high agreement level among LLMs, with an average $\rho=0.81$, despite their different providers. Conversely, they only moderately correlate with human experts, with an average $\rho=0.59$ and small variation across models.
@@
Political Biases
To probe political biases in LLM ratings, we focus on the 2,683 sources with partisan leanings in the US context. As mentioned in the Methods section, we query the LLMs with the same prompts except for assigning them different political identities. Specifically, for each source, we instruct the models to rate its credibility from the viewpoints of Democrats, Independents, and Republicans, respectively.
Figure 4: Heatmaps of Spearman correlation coefficients among the ratings generated by LLMs with different political identities. The highest correlation coefficients between the default LLM configuration and different political identities are highlighted by squares with solid edges. The highest correlation coefficients between human experts and different political identities are highlighted by squares with dashed edges.
In Figure 4, we show the correlations between the ratings generated by LLMs with different political identities and those from human experts. Ratings from LLMs in the default configuration (no political identities) are closest to those from the Independent identities, followed by Democratic identities. The correlations with the LLMs identified as Republicans are the lowest. When compared with human experts, the LLMs identified as Independents show higher correlations than all other cases, including the default ones. These patterns are consistent across all models.
To quantify the political biases of LLMs with different political identities, we calculate the LLM rating bias score for each source, defined as the difference between the LLM rating and the human expert rating. This metric accounts for the fact that left-leaning sources in our dataset tend to have higher human expert ratings. A positive/negative bias score means the LLM considers the source more/less credible than expected.
Figure 5: Distributions of LLM rating bias scores of GPT-4o mini with different political identities. The blue and red violins represent the results for left- and right-leaning sources, respectively. Significance of t-tests is indicated by ***: $p<0.001$, *:$p<0.05$, NS: not significant.
@@
In Figure 5, we show the distributions of LLM rating bias scores for GPT-4o mini with different political identities. We find that the default configuration and the Democratic identity exhibit liberal bias and are more likely to assign higher-than-expected credibility scores to left-leaning sources. The Republican identity, on the other hand, favors right-leaning sources. The Independent identity shows no significant differences in LLM rating bias scores between left- and right-leaning sources.
@@
Figure 6: Political biases of different LLM-identity configurations. The political biases are measured by t-statistics between the distributions of LLM rating bias scores on left- and right-leaning sources. A negative/positive t-statistic means the LLM-identity configuration favors left-/right-leaning news outlets. The solid symbols indicate statistically significant differences ($p<0.05$).
@@
We replicate the analysis for all LLMs and show the results in Figure 6. Due to space constraints, we only report the t-statistics from the tests comparing the distributions of LLM rating bias scores for left- and right-leaning news sources. For all eight models, assigning a Democratic identity leads to a liberal bias, whereas assigning a Republican identity leads to a conservative bias. The Independent identity shows no significant political biases for half of the models but a liberal bias for the other half, although less pronounced than the Democratic identity. Under the default configuration, where no explicit political identities are assigned to the LLMs, all models except for GPT-4 show a statistically significant liberal bias.
Political Bias and Rating Accuracy
@@
Figure 7: Political bias vs. rating accuracy for all LLM-identity configurations. We use the t-statistics from the tests comparing the distributions of LLM rating bias scores on left- and right-leaning sources to quantify the political bias, and the correlation with human experts for accuracy. LLM-identity configurations with left or right biases are separated and the lines represent linear regressions on the two groups.
@@
The results in Figures 4 and 6 suggest a negative correlation between the political biases and the accuracy of LLM-identity configurations. To confirm this, we present a scatter plot of all model-identity configurations in Figure 7. Regardless of the direction, we find that a stronger political bias correlates with lower rating accuracy. This finding suggests that the misalignment between LLMs and human experts is partially due to the political biases embedded in LLMs, and that reducing these biases could enhance rating accuracy.
@@
Pennycook and Rand (2019) ran an experiment in which participants produced news source ratings that correlated with their different ideologies. However, the combined ratings from participants on both sides correlate strongly with expert ratings. Given that LLMs exhibit consistent biases congruent with assigned identities, we test whether combining ratings from LLMs with different viewpoints could reduce overall biases and improve accuracy.
@@
We consider two aggregate ratings for each model: (1) D+R: the average ratings of the Democratic and Republican identities; and (2) D+I+R: the average ratings of the Democratic, Independent, and Republican identities. Although different weights can be assigned to these political identities, we choose equal weights to reflect the composition of American voters with varying partisanship.999pewresearch.org/politics/2024/04/09/the-partisanship-and-ideology-of-american-voters
@@
Figure 8: Accuracy of different LLM ratings. We show the Spearman correlation coefficients between different LLM ratings and those from human experts. For each model, we report the results for the default configuration, three political identities (Democrat, Independent, and Republican), as well as two aggregate ratings, D+R (average of Democratic and Republican identities) and D+I+R (average of all three political identities). The top three configurations for each model are annotated.
@@
In Figure 8, we compare the accuracy of D+R and D+I+R with the results for different political identities. In most cases, D+R and D+I+R ratings are more accurate than either Democratic or Republican ratings, suggesting that blending different perspectives leads to better judgment of source credibility. Across all configurations, the Independent identity is the winner for four models, while D+I+R leads in the other four cases. However, the differences between the Independent and D+I+R ratings are marginal. We, therefore, conjecture that aggregate ratings from partisan identities do not lead to significant improvement over the Independent identity, which already exhibits relatively low political bias.
Discussion
@@
In this paper, we systematically audit eight widely used LLMs from three major providers to test their ability to discern credible information sources from low-credibility ones. For most news sources tested in our experiments, the LLMs could provide ratings as instructed. However, larger models tend to refuse to rate less popular sources due to insufficient knowledge, while smaller models tend to hallucinate responses. Comparing the LLM ratings, these models exhibit a high level of agreement, despite being trained by different providers. Conversely, the models only moderately correlate with human experts.
@@
It remains unclear how the LLMs acquire this capacity. Based on the explanations provided by the models alongside the ratings, we hypothesize that the models summarize descriptions of the given news sources from their training data and generate ratings accordingly. For instance, when encountering high-credibility sources, the LLMs often note that these are well-established and reputable news websites. This could explain the high correlation among the LLMs, as they likely have common training data (Liu et al. 2024).
@@
Focusing on news sources with clear political leanings in the US context, we find that assigning partisan identities to LLMs constantly steers their ratings to favor sources with congruent leanings. This result indicates that LLMs can reflect the viewpoints of humans with different political ideologies (Argyle et al. 2023; Simmons 2023). In their default configurations, LLMs exhibit a liberal bias: they are more likely to assign higher-than-expected credibility ratings for left-leaning sources compared to right-leaning ones. LLMs identified as Independents are the least biased, but still show a weak liberal bias. These findings align with previous studies indicating that many LLMs have left-leaning tendencies (Santurkar et al. 2023).
Our auditing has some limitations. There are different approaches to prompt the LLMs, which might yield different outcomes. For instance, one could employ a binary classification approach or ask LLMs to rank two sources at a time (Qin et al. 2024). We were unable to test all of these approaches. Additionally, our study focuses on only eight representative models, a tiny portion of LLMs on the market. Given the rapid development in the field, new models with different behaviors will emerge soon.
Despite these limitations, we believe our findings provide valuable insights into the current state-of-the-art LLMs and could be generalized to other models considering that most results are consistent regardless of the model size and provider. Given that these LLMs have been deployed in widely used systems, the findings have important implications.
Our experiments show that LLMs have the potential to serve as an affordable and accessible reference for source credibility ratings. Such ratings are vital for researchers to study the dissemination and prevalence of misinformation online (Lazer et al. 2018; Lin et al. 2023). With the ability to provide contextual information and actively answer user questions, LLMs might also be adopted by the general public as scalable media literacy tools to investigate the credibility of news sources and perform fact-checking (Pennycook and Rand 2019; Hoes, Altay, and Bermeo 2023). On the one hand, the accessibility and low cost of LLMs make them an appealing alternative to human rating services. On the other hand, our results indicate that the accuracy of LLMs is not perfect, calling for further analysis and comparison with other misinformation intervention methods (Kozyreva et al. 2024). Additionally, LLMs may be vulnerable to manipulation. Bad actors could pollute LLM training data by creating fake news source review websites.
Our findings also highlight the challenges of using LLMs as information curators. Although not confirmed by the providers, leaked system prompts of popular LLMs suggest that they are configured to retrieve information according to the quality and credibility of sources when searching the Internet, along with other criteria such as the relevance to user questions and the freshness of the information.101010github.com/jujumilk3/leaked-system-prompts However, it is unclear how the LLMs understand and implement these criteria in production.
A famous failure story came from Google in 2024, when its AI overview feature produced bizarre answers such as recommending that people eat rocks.111111bbc.com/news/articles/cd11gzejgz4o According to Google, these odd responses originated from low-quality information sources,121212blog.google/products/search/ai-overviews-update-may-2024 yet their AI models appeared to overlook it. Although these errors may have a limited negative impact, they highlight the risks of using LLMs for information curation in high-stakes contexts such as public health and elections.
Our findings provide valuable insights about such risks. We show that LLMs may lack information about many information sources, especially unpopular ones. However, it is inevitable for LLMs to encounter unfamiliar sources when curating information for users, especially when they search for misinformation topics (Boyd and Golebiewski 2018; Aslett et al. 2024). Under these circumstances, forcing the models to produce summaries or responses could inadvertently amplify lesser-known low-credibility sources. Even when the models know the information sources, including the highly popular ones, their assessment of credibility still deviates from human judgment. This inaccuracy might suppress credible sources and amplify low-credibility ones. When dealing with political information, the inherent bias of the LLMs in the default configuration might lead to distorted outcomes. The finding that assigning partisan identities to LLMs could induce corresponding biases suggests that LLMs tailored to user preferences and views might exacerbate echo chambers and polarization.
@@
We also explore approaches to reduce political biases and increase the accuracy of the LLM ratings. Given the negative correlation between them, explicitly assigning an Independent identity to an LLM or mixing the ratings from different identities could help achieve both objectives simultaneously. However, even the unbiased ratings fail to align well with human judgment. Future models might consider explicitly integrating source credibility information into the training phases (Ouyang et al. 2022). When incorporating LLMs into larger systems such as search engines, one may also consider augmenting the models with external knowledge of the sources and instructing the models to give greater attention to credible ones.

### v3

@@ 4.2. Rating Accuracy
Figure 3. Heatmap of source credibility rating correlation (Spearman’s $\rho$) among different LLMs and human experts. Results in the upper right triangle of the heatmap are based on 3,077 (40.9%) sources rated by all LLMs. Results in the lower left triangle are based on the sources rated by both raters in comparison.
[…]
Heatmap of source credibility rating correlation (Spearman) among different LLMs and human experts.
Let us measure how well the LLM ratings correlate with each other and align with those from human experts. For each pair of raters (LLMs or human experts), we calculate the Spearman rank correlation coefficient $\rho$ between their ratings. Other metrics of performance, such as accuracy and F1 scores, are not used because they typically require a threshold to determine the positive/negative cases, which can be arbitrary in the present context. And because the distributions of the ratings are not normal (see visualizations in the Appendix), we use the Spearman rank correlation coefficient instead of the Pearson correlation coefficient.
@@
Given that some LLMs only provide ratings for a subset of the sources, we use two different approaches to calculate the correlation coefficients. First, we include only the 3,077 (40.9%) sources that are rated by all LLMs. Second, for each pair of raters, we focus on their intersection. The results from both approaches are shown in Figure 3 and demonstrate similar patterns. Therefore, our discussion below will focus on the results from the first approach.
@@
All correlation coefficients in Figure 3 are positive and statistically significant with $p<0.001$. We find a high agreement level among LLMs, with an average $\rho=0.79$, despite their different providers. Conversely, they only moderately correlate with human experts, with an average $\rho=0.50$ and small variation across models.
To provide context for our findings, we compare them with the recent work of Williams et al. (2024). Their study combines social media data and web graph analysis to identify unreliable domains, achieving slightly better performance (accuracy=0.819) than GPT-3.5 (accuracy=0.782) on the same set of domains used in our study. While a direct comparison is not possible due to the unavailability of their code and data, we test GPT-3.5 (gpt-3.5-turbo-0125) under our experimental conditions. The model yields a Spearman correlation coefficient of 0.50 with human expert ratings, which is the same as the average correlation coefficient across all LLMs in our study.
@@
4.3. Political Biases
To probe political biases in LLM ratings, we focus on the 2,683 sources with partisan leanings in the US context. We query the LLMs with the same prompts except for assigning them different partisan roles. Specifically, for each source, we instruct the models to rate its credibility from the viewpoints of Democrats, Independents, and Republicans, respectively.
Figure 4. Heatmaps of Spearman correlation coefficients among the ratings generated by LLMs with different partisan roles. The highest correlation coefficients between the default LLM configuration and different partisan roles are highlighted by squares with solid edges. The highest correlation coefficients between human experts and different partisan roles are highlighted by squares with dashed edges.
[…]
Heatmaps of Spearman correlation coefficients among the ratings generated by LLMs with different partisan roles.
In Figure 4, we show the correlations between the ratings generated by LLMs with different partisan roles and those from human experts. Ratings from LLMs in the default configuration (no partisan roles) are closest to those from the Independent role, followed by Democratic roles. The correlations with the LLMs identified as Republicans are the lowest. When compared with human experts, the LLMs identified as Independents show higher correlations than all other cases, including the default ones. These patterns are consistent across all models.
To quantify the partisan biases of LLMs with different partisan roles, we calculate the LLM rating bias score for each source, defined as the difference between the LLM rating and the human expert rating. This metric accounts for the fact that left-leaning sources in our dataset tend to have higher human expert ratings. A positive/negative bias score means the LLM considers the source more/less credible than expected.
Figure 5. Distributions of LLM rating bias scores of GPT-4o mini with different partisan roles. The blue and red violins represent the results for left- and right-leaning sources, respectively. Significance of t-tests is indicated by ***: $p<0.001$, *:$p<0.05$, NS: not significant.
[…]
Distributions of LLM rating bias scores of GPT-4o mini with different partisan roles.
@@ 4.3. Political Biases
In Figure 5, we show the distributions of LLM rating bias scores for GPT-4o mini with different partisan roles. We find that the default configuration and the Democratic role exhibit liberal bias and are more likely to assign higher-than-expected credibility scores to left-leaning sources. The Republican role, on the other hand, favors right-leaning sources. The Independent role shows no significant differences in LLM rating bias scores between left- and right-leaning sources.
@@
Figure 6. Political biases of different LLM-role configurations. The political biases are measured by t-statistics between the distributions of LLM rating bias scores on the left- and right-leaning sources. A negative/positive t-statistic means the LLM-role configuration favors left-/right-leaning news outlets. The solid symbols indicate statistically significant differences ($p<0.05$).
[…]
Political biases of different LLM-role configurations.
@@
We replicate the analysis for all LLMs and show the results in Figure 6. Due to space constraints, we only report the t-statistics from the tests comparing the distributions of LLM rating bias scores for left- and right-leaning news sources. We find that assigning a Democratic role leads to a liberal bias for all models except for o1-mini, whereas assigning a Republican role leads to a conservative bias for all models. The Independent role shows no significant political biases for five models but a liberal bias for the other four models, although less pronounced than the Democratic role. Under the default configuration, where no explicit partisan roles are assigned to the LLMs, all models except for GPT-4 and o1-mini show a statistically significant liberal bias.
4.4. Political Bias and Rating Accuracy
@@
Figure 7. Political bias vs. rating accuracy for all LLM-role configurations. We use the t-statistics from the tests comparing the distributions of LLM rating bias scores on left- and right-leaning sources to quantify the political bias and the correlation with human experts for accuracy. LLM-role configurations with left or right biases are separated and the lines represent linear regressions on the two groups. For the left-biased data points, we have the Spearman correlation coefficient between the political bias and the rating accuracy of $\rho=0.18$ ($p=0.38$). For the right-biased data points, we have $\rho=-0.67$ ($p=0.02$).
[…]
Political bias vs. rating accuracy for all LLM-role configurations.
@@
The results in Figures 4 and 6 suggest a negative correlation between the political biases and the accuracy of LLM-role configurations. To confirm this, we present a scatter plot of all model-role configurations in Figure 7. Regardless of the direction, we find that a stronger political bias correlates with lower rating accuracy, although such correlation is only statistically significant for the cases where the LLM-role configuration is right-biased. This finding suggests that the misalignment between LLMs and human experts is partially due to the political biases embedded in LLMs and that reducing these biases could enhance rating accuracy.
@@
Pennycook and Rand (2019) ran an experiment in which participants produced news source ratings that correlated with their different ideologies. However, the combined ratings from participants on both sides correlate strongly with expert ratings. Given that LLMs exhibit consistent biases congruent with assigned roles, we test whether combining ratings from LLMs with different viewpoints could reduce overall biases and improve accuracy.
@@
We consider two aggregate ratings for each model: (1) D+R, the average ratings of the Democratic and Republican roles, and (2) D+I+R, the average ratings of the Democratic, Independent, and Republican roles. Although different weights can be assigned to these partisan roles, we choose equal weights to reflect the composition of American voters with varying partisanship.121212pewresearch.org/politics/2024/04/09/the-partisanship-and-ideology-of-american-voters
@@
Figure 8. Accuracy of different LLM ratings. We show the Spearman correlation coefficients between different LLM ratings and those from human experts. For each model, we report the results for the default configuration, three partisan roles (Democrat, Independent, and Republican), as well as two aggregate ratings, D+R (average of Democratic and Republican roles) and D+I+R (average of all three political roles). The top three configurations are annotated.
[…]
Accuracy of different LLM ratings.
@@
In Figure 8, we compare the accuracy of D+R and D+I+R with the results for different partisan roles. In most cases, D+R and D+I+R ratings are more accurate than either Democratic or Republican ratings, suggesting that blending different perspectives leads to better judgment of source credibility. Across all configurations, the Independent role is the winner for four models, while D+I+R leads in the other four cases. However, the differences between the Independent and D+I+R ratings are marginal. We, therefore, conjecture that aggregate ratings from partisan roles do not lead to significant improvement over the Independent role, which already exhibits relatively low political bias.
5. Discussion
5.1. Summary of Results
@@
In this paper, we systematically audit nine widely used LLMs from three leading providers to test their ability to discern credible information sources from low-credibility ones. For most news sources tested in our experiments, the LLMs could provide ratings as instructed. However, larger models tend to refuse to rate less popular sources due to insufficient knowledge, while smaller models tend to make more errors. Comparing the LLM ratings, these models exhibit a high level of agreement despite being trained by different providers. Conversely, the models only moderately correlate with human experts.
@@
It remains unclear how the LLMs acquire this capacity. Based on the explanations provided by the models alongside the ratings, we hypothesize that the models summarize descriptions of the given news sources from their training data and generate ratings accordingly. For instance, when encountering high-credibility sources, the LLMs often note that these are well-established and reputable news websites. This could explain the high correlation among the LLMs, as they likely have common training data (Liu et al., 2024).
@@
Focusing on news sources with clear political leanings in the US context, we find that assigning partisan roles to LLMs steers their ratings to favor sources with congruent leanings in most cases. This result indicates that LLMs can reflect the viewpoints of humans with different political ideologies (Argyle et al., 2023; Simmons, 2023). In their default configurations, most LLMs exhibit a liberal bias: they are more likely to assign higher-than-expected credibility ratings for left-leaning sources compared to right-leaning ones. LLMs assigned with an Independent role are the least biased but still show a weak liberal bias. These findings align with previous studies indicating that many LLMs have left-leaning tendencies (Santurkar et al., 2023).
@@
We also explore approaches to reduce political biases and increase the accuracy of the LLM ratings. We find that explicitly assigning an Independent role to an LLM or mixing the ratings from different roles could help achieve both objectives simultaneously. However, even unbiased ratings fail to align perfectly with human judgment.
5.2. Limitations
Our auditing has some limitations. First, since the evaluation relies on existing human ratings, any biases in these ratings may propagate into our findings. For example, our binary classification of news sources as left- or right-leaning based on a zero threshold applied to the scores by Robertson et al. (2018) simplifies our bias analysis, but the alternative 5-class approach used by organizations like MBFC might provide more nuanced results.
In our experiments, we only provide the LLMs with the domains of the news sources, which is useful for assessing the completeness and accuracy of the internal knowledge. However, in real-world scenarios, the models are likely to have access to additional information about the sources, such as the metadata and content of the pages, which could potentially improve their performance (Schlichtkrull, 2024).
There are different approaches to prompt the LLMs, which might yield different outcomes. For instance, one could employ a binary classification approach or ask LLMs to rank two sources at a time (Qin et al., 2024). When assigning a partisan role to the LLMs, one could also provide them with detailed demographics and background information about a persona. Additionally, different prompt engineering techniques could be deployed to boost the performance of the LLMs. We were unable to test all of these approaches.
Our findings might not be generalizable to all LLMs and contexts. Despite our efforts to test as many representative models as possible, we are unable to cover all the other LLMs from different providers on the market. Given the rapid development in the field, new models with different behaviors will emerge soon. Finally, our experiments mainly focus on the US context, which might not be representative of other countries and non-English news.
5.3. Implications
Despite these limitations, we believe our findings provide valuable insights into the current state-of-the-art LLMs, considering that the general patterns are consistent regardless of the model size and provider. Given that these LLMs have been deployed in widely used systems, the findings have important implications.
Our experiments show that LLMs have the potential to serve as an affordable and accessible reference for source credibility ratings. Such ratings are vital for researchers to study the dissemination and prevalence of misinformation online (Lazer et al., 2018; Lin et al., 2023). With the ability to provide contextual information and actively answer user questions, LLMs might also be adopted by the general public as scalable media literacy tools to investigate the credibility of news sources and perform fact-checking (Pennycook and Rand, 2019; Hoes et al., 2023). However, our results indicate that the accuracy of LLMs is not perfect, calling for further analysis and comparison with other misinformation intervention methods (Kozyreva et al., 2024). Additionally, LLMs may be vulnerable to manipulation. For instance, bad actors could pollute LLM training data by creating fake news source review websites.
Our findings also highlight the challenges of using LLMs as information curators. Although not confirmed by the providers, leaked system prompts of popular LLMs suggest that they are configured to retrieve information according to the quality and credibility of sources when searching the Internet, along with other criteria such as the relevance to user questions and the freshness of the information.131313github.com/jujumilk3/leaked-system-prompts However, it is unclear how the LLMs understand and implement these criteria in production.
A famous failure story came from Google in 2024 when its AI overview feature produced bizarre answers such as recommending that people eat rocks.141414bbc.com/news/articles/cd11gzejgz4o According to Google, these odd responses originated from low-quality information sources,151515blog.google/products/search/ai-overviews-update-may-2024 yet their AI models appeared to overlook it. Although these errors may have a limited negative impact, they highlight the risks of using LLMs for information curation in high-stakes contexts such as public health and elections.
Our findings provide valuable insights about such risks. We show that LLMs may lack information about many information sources, especially unpopular ones. However, it is inevitable for LLMs to encounter unfamiliar sources when curating information for users, especially when they search for misinformation topics (Boyd and Golebiewski, 2018; Aslett et al., 2024). Under these circumstances, forcing the models to produce summaries or responses could inadvertently amplify lesser-known low-credibility sources. Even when the models know the information sources, including the highly popular ones, their assessment of credibility still deviates from human judgment. This inaccuracy might suppress credible sources and amplify low-credibility ones. When dealing with political information, the inherent bias of the LLMs in the default configuration might lead to distorted outcomes. The finding that assigning partisan roles to LLMs could induce corresponding biases suggests that LLMs tailored to user preferences and views might exacerbate echo chambers and polarization.
5.4. Future Studies
For future research, a critical challenge lies in mitigating biases and enhancing the accuracy of LLMs in assessing source credibility. While one might expect that models demonstrating superior performance on standard benchmarks would naturally excel at this specific task, our findings suggest otherwise. Our comparative analysis of models within the same family, such as the GPT series, reveals that increased model size and parameter count do not consistently translate to improved accuracy or reduced biases. Although OpenAI’s o1 model series represents an innovative approach by training models specifically for complex reasoning tasks, our evaluation of o1-mini indicates that this enhanced reasoning capability has not effectively transferred to the present task. These findings suggest that more capable LLMs in the future might not necessarily perform better on the source credibility assessment task.
In our attempt to address this challenge, we leverage the prompt-based approach by assigning partisan roles to the LLMs and aggregating ratings from multiple roles. This approach is simple and cheap, and can be applied to both proprietary and open-source models, but the improvement is limited. Future models might consider explicitly integrating source credibility information into the models either through fine-tuning or including the data in the pre-training phase (Ouyang et al., 2022). When embedding LLMs into larger systems such as search engines, one may also consider providing the models with the credibility ratings of the sources together with the content and query for reasoning.
Another key challenge is to enhance LLMs’ ability to evaluate previously unseen information sources. Our findings demonstrate that LLMs struggle to assess the credibility of less popular sources, significantly limiting their effectiveness as information curators. While human expert evaluation could provide accurate ratings, this approach is neither sustainable nor scalable, given the vast number of sources and limited expert resources. Future research should explore automated frameworks that leverage LLMs to assess source credibility (Schlichtkrull, 2024). Such frameworks could autonomously gather information on the sources from the Internet, leveraging advanced reasoning capabilities to process such information, and generate reliable credibility ratings.